The Velocity-Time graphs of the two cars; K and L are given
below together. If the cars start their journey together and they
have a position difference of 10 m at time t, what will be the
position difference between them at time 4t at the end?
1
Expert's answer
2020-05-08T16:12:40-0400
"s_k-s_l=0.5(t+3t)(2v)-0.5(t+3t)(2v)+10\\=0+10=10\ m"